# The StudyChat Dataset: Exploring Student Dialogues With ChatGPT in an Artificial Intelligence Course

### v3

@@ Abstract
The widespread availability of large language models (LLMs), such as ChatGPT, has significantly impacted education, raising both opportunities and challenges. Students can frequently interact with LLM-powered, interactive learning tools, but their usage patterns need to be monitored and understood. We introduce StudyChat, a publicly available dataset capturing real-world student interactions with an LLM-powered tutoring chatbot in a semester-long, university-level artificial intelligence (AI) course. We deploy a web application that replicates ChatGPT’s core functionalities, and use it to log student interactions with the LLM while working on programming assignments. We collect 16,851 interactions, which we annotate using a dialogue act labeling schema inspired by observed interaction patterns and prior research. We analyze these interactions, highlight usage trends, and analyze how specific student behavior correlates with their course outcome. We find that students who prompt LLMs for conceptual understanding and coding help tend to perform better on assignments and exams. Moreover, students who use LLMs to write reports and circumvent assignment learning objectives have lower outcomes on exams than others. StudyChat serves as a shared resource to facilitate further research on the evolving role of LLMs in education.  [^1] https://huggingface.co/datasets/wmcnicho/StudyChat.
Keywords Conversational Datasets  $\cdot$ CS Education  $\cdot$ Dialogue Analysis  $\cdot$ Large Language Models  $\cdot$ Natural Language Processing
1 Introduction
@@
Advances in large language models (LLMs) have had major ramifications on almost all levels of education. Many LLM-powered educational technologies have recently been developed, leveraging the generative capabilities of these models to create new course content (MacNeil et al., 2022), assignments (MacNeil et al., 2023), and interactive learning experiences (Jin et al., 2024; Taneja et al., 2024). These interactive technologies enable rich engagement with course material by allowing students to engage in on-demand conversations with an interactive tutor, and receive personalized feedback on their work. However, while the potential use cases for LLMs in education are vast, many educators are understandably concerned about student misuse of LLM chatbot tools, such as ChatGPT Lau and Guo (2023). While a motivated learner can use these tools positively, an unmotivated student could use them to complete their assignments, circumventing the learning objectives of the assignment, or become over-reliant on the model (Bastani et al., 2024).
@@
This dichotomy between learning opportunity and academic risk necessitates an exploration into understanding student behavior with LLM chatbot tools. A more nuanced understanding of this behavior could inform further development of LLM-based Intelligent Tutoring Systems (ITS) that dynamically adapt to students, encouraging learning-focused behaviors and discouraging counterproductive ones. Furthermore, a deeper understanding of student behavior can lead to more robust student models, which better estimate knowledge states, misconceptions, and learning trajectories. For example, such a model could potentially detect when a student is overly reliant on LLM responses and flag this behavior to instructors in real-time, or be used to help develop assessments which are specifically designed to test for this over-reliance.
@@
In this work, we conduct a year-long case study to explore student behavior with an LLM, specifically in the context of upper-division, university-level Computer Science (CS) education. We develop a web application with the same core features and underlying LLM as ChatGPT, and deploy it in an introductory-level course on Artificial Intelligence (AI). We encourage students to use this tool, without restrictions, on programming assignments over the course of the semester. Across two semesters and with student consent we recorded 2,214 student-LLM conversations from 203 students across 7 unique assignments, totaling 16,851 unique utterances. We also collect 924 programming assignment submissions from 158 students who further agreed to share their work. In addition, we annotate these conversation utterances with a dialogue act (DA) labeling schema; the schema is inspired by existing studies on co-programming behavior and our systematic review of the conversations. We leverage an LLM prompting-based approach to do this annotation at scale and validate this approach with a human evaluation.
@@
We explore how students engage with an LLM during programming assignments. In doing so our primary research objective is to uncover patterns of behavior across students and analyze if these trends correlate with course outcomes. To do this, we analyze these DA labels by performing regression and clustering analysis. Our key findings are: 1) students primarily use the LLM to ask questions, but our findings suggest that some question types are negatively correlated with outcome, suggesting student confusion possibly due to LLM hallucination; 2) certain behavior patterns such as asking conceptual questions or help in code writing are positively associated with both assignment and held out exam scores; 3) students with low and high usage of LLMs have similar average outcome but students that use that LLMs with high frequency tend to have less variance in a more narrow positive range; 4) clustering analysis reveals students who use the LLM primarily for code writing and not other writing requests tend to score higher on both assignments and exams. These findings provide data-driven insights into patterns of student LLM behaviors in a university-level CS course. We release this dataset to the broader research community to foster collaboration and exploration into LLMs’ impact on AI and CS education.
2 Related Works
2.1 ChatGPT Classroom Studies
@@
Since its release in November 2022, there has been a vast amount of research interest into how ChatGPT can be utilized in classroom settings. Studies have explored various generative applications of ChatGPT for course exercises (Xiao et al., 2023), code explanations (Chen et al., 2023), multiple-choice distractors (Feng et al., 2024), and more (Khan et al., 2023; Li et al., 2024; Hsieh and Yang, 2024). Direct classroom studies have also explored how ChatGPT-powered applications can answer student questions as a virtual teaching assistant (Taneja et al., 2024), help students learn by teaching chatbots to code (Jin et al., 2024), and even provide coaching and feedback to instructors on content delivery (Wang and Demszky, 2023).
@@
Closely related to this work are a handful of cases studies exploring student use of ChatGPT in university level CS classrooms. Ghimire et. al explore student use of ChatGPT for two assignments in a CS1 programming course, finding students often ask questions of the system instead of directly seeking solutions (Ghimire and Edwards, 2024). Ma et. al also perform a case study where CS1 students were allowed to use ChatGPT freely on in-class exercises and on a course project, but only for conceptual questions on homework citepma2024enhancing. Additionally, Brender et. al perform a case study on how graduate students in a robotics course utilize ChatGPT on a voluntary assignment, controlling for learning outcomes and clustering behavior into broad categories (Brender et al., 2024). Analysis of student prompts in these studies have noted a student trend to use models to ask questions more than directly writing code solutions which is consistent with our observations and those by Bhalerao (2024). Surveys on student perceptions towards LLMs have also been conducted which suggest that CS student adoption of LLMs is nuanced and not universal (Weber et al., 2024). Similarly, we observe that a some students use the system heavily while most students engage moderately with the model, revealing a range of adoption behaviors.
2.2 Education Discourse Analysis
@@
It has been demonstrated by many that tutoring is a highly effective form of instruction Cohen et al. (1982). Thus, human tutoring dialogue has been studied extensively Graesser et al. (1995). In recent years, DA classification on conversation turns is often used in this discourse analysis Chen et al. (2011). DA labels classify a participant’s utterance by the function it serves in the conversation. When DA labeling is applied to an entire corpus of dialogues, it enables discourse-level analysis of conversation patterns that are frequent or significant. In education, significant patterns are often those associated with better learning outcomes, as such existing works have studied such patterns in tutor-student or student-student learning environments Graesser et al. (1995); Earle-Randell et al. (2023). Existing works also analyze human-machine classroom conversations between students and ITS chatbots Vail and Boyer (2014); Graesser et al. (2004); Fisher et al. (2020). These works provide promising insights which can inform more effective tutoring pedagogy Rus et al. (2017) and interactive learning tools Chen et al. (2011).
@@
While DA analysis is an attractive approach to discourse analysis, the labeling process poses many challenges, solutions to which constitute an large body of work Boyer et al. (2011); Tan et al. (2023); Ezen-Can and Boyer (2015); Lin et al. (2023); Boyer et al. (2010); Kumaran et al. (2023). First, there exists the task of annotating a large corpus of conversations. To automate such a task, various machine learning based approaches to this have been used such as Hidden Markov Models Boyer et al. (2010), decision trees Samei et al. (2014), and Markov Random Fields Ezen-Can and Boyer (2015). In our work, we employ a LLM prompt-based approach to DA labeling, similar to that of Wang et al. (2023). Second, there exists the challenge of robust schema design, where existing work has looked into refining prior labeling schemata to be less ambiguous Vail and Boyer (2014) and to try to develop classroom agnostic labeling schema Hennessy et al. (2016).
2.3 Education Dialogue Datasets
@@
In an effort to collectively improve data-driven learning technologies, there have been recent efforts to expand the amount of education dialogue datasets in a variety of learning domains. Datasets such as TalkMoves and NCTE directly capture and annotate classroom discourse in K-12 lessons Suresh et al. (2022); Demszky and Hill (2022). Other works employ crowd workers or leverage LLMs to simulate the student or teacher role, such as MathDial Macina et al. (2023), Teacher-Student Chatroom Corpus (TSCC) Caines et al. (2020), Conversational Instruction with Multi-responses and Actions (CIMA) Stasaski et al. (2020), and GrounDialog Zhang et al. (2023). Our work contributes to this growing body of research by providing a large-scale, annotated dialogue dataset of in a university-level CS course, offering insights into how students engage with LLM-chatbot tools in a real learning environment.
Figure 1: The web application used in this study. The user interface mirrors functionality of ChatGPT and allows students to organize conversations by assignment.
3 Dataset
Table 1: Dialogue act schema for student interactions with ChatGPT.
Category Acts Description
Writing Write Code, Write English, Conversion, Summarize, Other The user asks the LLM to generate text, such as code, explanations, or summaries.
Editing Edit Code, Edit English, Other The user asks for modifications or improvements to existing code or text.
Contextual Questions Assignment Clarification, Code Explanation, Interpret Output, Other The user asks a question based on something from the assignment or prior conversation.
Conceptual Questions Programming Language, Python Library, Computer Science, Programming Tools, Mathematics, Other Concept The user asks a general question about programming, libraries, tools, or CS concepts unrelated to the conversation.
Verification Verify Code, Verify Report, Verify Output, Other The user requests validation of their code, writing, or output to ensure correctness.
Context Assignment Information, Error Message, Code, Other The user supplies relevant information to the LLM, such as assignment instructions, code snippets, or error messages.
Off Topic Chit-Chat, Greeting, Gratitude, Other Conversations unrelated to the assignment, such as casual discussions, greetings, or expressions of gratitude.
Misc Other A catch-all category for dialogue acts that do not fit elsewhere.
@@ 3 Dataset
3.1 Participants and Course Context
Our IRB-approved study was conducted among students aged 18 or above, who were enrolled in an AI course at a US R1 university in the Fall 2024 and Spring 2025 semesters. Across both semesters, 295 students were enrolled in the course, of which 203 agreed to participate in the study. Students who did not participate were also given equivalent access to use the tool over the semester, but their data was not collected. Students, were in their 3rd/4th years, had completed their foundational Computer Science course sequences and were taking higher-level elective courses.
@@ 3.1 Participants and Course Context
The programming assignments are Python-based and can be broadly categorized into two types. The first assignment type is a tutorial, where students are given incomplete Jupyter notebooks which guide students to learn fundamentals operations in applied machine learning Python libraries including Pandas, SciKit Learn, and PyTorch. The second assignment type is an applied scenario, where students are given Python starter code which they extend by applying an algorithm or concept taught in class. For example, in assignment 2, students are tasked with creating a text autocomplete engine by implementing various search algorithms to traverse graphs to find word completions. Assignments were altered slightly between semesters, but generally cover the same topics. All assignment instruction details, starter code, and 924 graded submissions are included in the dataset.
3.2 Data Collection
@@
In order to effectively collect student utterances, we developed and deployed a web application designed to function similar to consumer LLM chatbot interfaces, shown in Figure 1. As our backing chatbot LLM, we used gpt-4o-mini, the LLM for the free tier of ChatGPT at the time of this study gpt . We evaluated this model by testing its capabilities with the libraries and concepts covered in the assignments, and found the model generally provides accurate guidance on the assignments and implementations that are often accurate. Students were instructed they could use this ChatGPT-based tool, without restrictions, on all assignments in the course. It was emphasized that their approach to using the tool would not affect their grade and they were encouraged to use the tool in any manner of their choosing. We used the default "You are a helpful assistant" system prompt, which does not discouraging answer-seeking behavior, and did not provide the chatbot with any course materials for reference. Our goal was to encourage students to use our platform instead of seeking an outside application for assignment help, and that the collected interactions would more accurately reflect how students generally use non-instructional, consumer chatbots.
3.3 Filtering of Student Identifiable Information
@@
To remove any Personally Identifiable Information (PII) from the conversations and assignment submissions, we systematically analyze the conversations to uncover patterns of PII exposure. We first randomly sample 12 conversations from each course topic and read through the conversations, scanning for traces of PII. Our main finding of potentially sensitive student information is in the form of pasted terminal outputs and error messages, which sometimes contain PII, such as a user’s directory or GitHub account name. Compiling these common patterns, we then develop a regular expression-based script to remove these patterns and any direct matches to known user PII in the course roster. Across the entire dataset, this process removes 6,413 cases of possible PII exposure. Then, we sample 12 additional conversations from each topic and performed the same procedure. In our second pass of the second set of conversations, we did not find any additional patterns or cases of PII exposure. We acknowledge that although seemingly effective, this process does not guarantee the removal of all PII exposures.
3.4 Data Statistics

### v4

# The StudyChat Dataset: Analyzing Student Dialogues With ChatGPT in an Artificial Intelligence Course

@@ Abstract.
The widespread availability of large language models (LLMs), such as ChatGPT, has significantly impacted education, raising both opportunities and challenges. Students can frequently interact with LLM-powered, interactive learning tools, but their usage patterns need to be observed and understood. We introduce StudyChat, a publicly available dataset capturing real-world student interactions with an LLM-powered tutoring chatbot in a semester-long, university-level artificial intelligence (AI) course. We deploy a web application that replicates ChatGPT’s core functionalities, and use it to log student interactions with the LLM while working on programming assignments. We collect 16,851 interactions, which we annotate using a dialogue act labeling schema inspired by observed interaction patterns and prior research. We analyze these interactions, highlight usage trends, and analyze how specific student behavior correlates with their course outcome. We find that students who prompt LLMs for conceptual understanding and coding help tend to perform better on assignments and exams. Moreover, students who use LLMs to write reports and circumvent assignment learning objectives have lower outcomes on exams than others. StudyChat serves as a shared resource to facilitate further research on the evolving role of LLMs in education  [^1] https://huggingface.co/datasets/wmcnicho/StudyChat.
Large Language Models, Educational Technology, Dialogue Acts, Student Behavior, Computer Science Education, ChatGPT, Dataset
journalyear: 2026copyright: noneconference: LAK26: 16th International Learning Analytics and Knowledge Conference; April 27-May 01, 2026; Bergen, Norwaybooktitle: LAK26: 16th International Learning Analytics and Knowledge Conference (LAK 2026), April 27-May 01, 2026, Bergen, Norwaydoi: 10.1145/3785022.3785029isbn: 979-8-4007-2066-6/2026/04ccs: Applied computing Interactive learning environmentsccs: Applied computing Computer-assisted instructionccs: Computing methodologies Discourse, dialogue and pragmatics
1. Introduction
@@
Advances in large language models (LLMs) have had major ramifications on almost all levels of education. Many LLM-powered educational technologies have recently been developed, leveraging the generative capabilities of these models to create new course content (MacNeil et al., 2022), assignments (MacNeil et al., 2023), and interactive learning experiences (Jin et al., 2024; Taneja et al., 2024). These interactive technologies enable rich engagement with course material by allowing students to engage in on-demand conversations with an interactive tutor and receive personalized feedback on their work. However, while the potential use cases for LLMs in education are vast, many educators are understandably concerned about student misuse of LLM chatbot tools, such as ChatGPT (Lau and Guo, 2023). While a motivated learner can use these tools positively, an unmotivated student could use them to complete their assignments, circumventing the learning objectives of the assignment, or become over-reliant on the model (Bastani et al., 2024).
@@
This dichotomy between learning opportunity and academic risk necessitates an exploration into understanding student behavior with LLM chatbot tools. A more nuanced understanding of this behavior could inform further development of LLM-based Intelligent Tutoring Systems (ITS) that dynamically adapt to students, encouraging learning-focused behaviors and discouraging counterproductive ones. Furthermore, a deeper understanding of student behavior can lead to more robust student models, which better estimate knowledge states, misconceptions, and learning trajectories. For example, such a model could potentially detect when a student is overly reliant on LLM responses and flag this behavior to instructors in real-time, or be used to help develop assessments that are specifically designed to test for this over-reliance.
Figure 1. An overview of the dataset collection and analysis pipeline present in our work.
An overview of the dataset collection and analysis pipeline present in this work.
@@
In this work, we conduct a year-long case study to explore student behavior with an LLM, specifically in the context of upper-division, university-level Computer Science (CS) education. We develop a web application with the same core features and underlying LLM as ChatGPT, and deploy it in an introductory-level course on Artificial Intelligence (AI). We encourage students to use this tool, without restrictions, on programming assignments over the course of the semester. Across two semesters and with student consent, we recorded 2,214 student-LLM conversations from 203 students across 7 unique assignments, totaling 16,851 unique utterances. We also collect 924 programming assignment submissions from 158 students who further agreed to share their work. In addition, we annotate these conversation utterances with a dialogue act (DA) labeling schema; the schema is inspired by existing studies on co-programming behavior and our systematic review of the conversations. We leverage an LLM prompting-based approach to do this annotation at scale and validate this approach with a human evaluation.
@@
We explore how students engage with an LLM during programming assignments. In doing so our primary research objective is to uncover patterns of behavior across students and analyze if these trends correlate with course outcomes. To do this, we analyze these DA labels by performing regression and clustering analysis. Our key findings are: 1) students primarily use the LLM to ask questions, but our findings suggest that some question types are negatively correlated with outcome, suggesting student confusion possibly due to LLM hallucination; 2) certain behavior patterns such as asking conceptual questions or help in code writing are positively associated with both assignment and held out exam scores; 3) students with low and high usage of LLMs have similar average course outcome but students that use that LLMs with high frequency tend to have less variance in a more narrow positive range; 4) clustering analysis reveals students who use the LLM primarily for code writing and not other writing requests tend to score higher on both assignments and exams. These findings provide data-driven insights into patterns of student LLM behaviors in a university-level CS course. We release this dataset to the broader research community to foster collaboration and exploration into LLMs’ impact on AI and CS education.
2. Related Works
2.1. ChatGPT Classroom Studies
@@
Since its release in November 2022, there has been a vast amount of research interest into how ChatGPT can be utilized in classroom settings (Zheng et al., 2025; Ammari et al., 2025). Studies have explored various applications of ChatGPT for course exercises (Xiao et al., 2023; Kumar et al., 2023), code explanations (Chen et al., 2023), multiple-choice distractors (Feng et al., 2024; Parikh et al., 2025), at-risk student detection (Liu et al., 2024), and more (Khan et al., 2023; Li et al., 2024; Hsieh and Yang, 2024; Dunder et al., 2024; Hedlin et al., 2025; Moore et al., 2022; Lee et al., 2024, 2025; Duan et al., 2025). Direct classroom studies have also explored how ChatGPT-powered applications can answer student questions as a virtual teaching assistant (Taneja et al., 2024), help students learn by teaching chatbots to code (Jin et al., 2024), and even provide coaching and feedback to instructors on content delivery (Wang and Demszky, 2023).
@@
Closely related to this work are a handful of case studies exploring student use of ChatGPT in university-level CS classrooms. Ghimire et. al explore student use of ChatGPT for two assignments in a CS1 programming course, finding students often ask questions of the system instead of directly seeking solutions (Ghimire and Edwards, 2024). Ma et. al also perform a case study where CS1 students were allowed to use ChatGPT freely on in-class exercises and on a course project, but only for conceptual questions on homework (Ma et al., 2024). Additionally, Brender et. al perform a case study on how graduate students in a robotics course utilize ChatGPT on a voluntary assignment, controlling for learning outcomes and clustering behavior into broad categories (Brender et al., 2024). Analysis of student prompts in these studies have noted a student trend to use models to ask questions more than directly writing code solutions, which is consistent with our observations and those by (Bhalerao, 2024). Surveys on student perceptions towards LLMs have also been conducted, which suggest that CS student adoption of LLMs is nuanced and not universal (Weber et al., 2024). Similarly, we observe that some students use the system heavily while most students engage moderately with the model, revealing a range of adoption behaviors.
2.2. Education Discourse Analysis
@@
It has been demonstrated by many that tutoring is a highly effective form of instruction (Cohen et al., 1982). Thus, human tutoring dialogue has been studied extensively (Graesser et al., 1995). In recent years, DA classification on conversation turns is often used in this discourse analysis (Chen et al., 2011; Ezen-Can et al., 2015). DA labels classify a participant’s utterance by the function it serves in the conversation. When DA labeling is applied to an entire corpus of dialogues, it enables discourse-level analysis of conversation patterns that are frequent or significant. In education, significant patterns are often those associated with better learning outcomes, as such existing works have studied such patterns in tutor-student or student-student learning environments (Graesser et al., 1995; Earle-Randell et al., 2023; Ikram et al., ). Existing works also analyze human-machine classroom conversations between students and ITS chatbots (Vail and Boyer, 2014; Graesser et al., 2004; Fisher et al., 2020). These works provide promising insights which can inform more effective tutoring pedagogy (Rus et al., 2017) and interactive learning tools (Chen et al., 2011).
@@
While DA analysis is an attractive approach to discourse analysis, the labeling process poses many challenges, solutions to which constitute a large body of work (Boyer et al., 2011; Tan et al., 2023; Ezen-Can and Boyer, 2015; Lin et al., 2023; Boyer et al., 2010; Kumaran et al., 2023). First, there exists the task of annotating a large corpus of conversations. To automate such a task, various machine learning based approaches to this have been used, such as Hidden Markov Models (Boyer et al., 2010), decision trees (Samei et al., 2014), and Markov Random Fields (Ezen-Can and Boyer, 2015). In our work, we employ a LLM prompt-based approach to DA labeling, similar to that of (Wang et al., 2023). Second, there exists the challenge of robust schema design, where existing work has looked into refining prior labeling schemata to be less ambiguous (Vail and Boyer, 2014) and to try to develop a classroom agnostic labeling schema (Hennessy et al., 2016).
2.3. Education Dialogue Datasets
@@
In an effort to collectively improve data-driven learning technologies, there have been recent efforts to expand the amount of education dialogue datasets in a variety of learning domains. Datasets such as TalkMoves and NCTE directly capture and annotate classroom discourse in K-12 lessons (Suresh et al., 2022; Demszky and Hill, 2022). Other works employ crowd workers or leverage LLMs to simulate the student or teacher role, such as MathDial (Macina et al., 2023), Teacher-Student Chatroom Corpus (TSCC) (Caines et al., 2020), Conversational Instruction with Multi-responses and Actions (CIMA) (Stasaski et al., 2020), and GrounDialog (Zhang et al., 2023). Our work contributes to this growing body of research by providing a large-scale, annotated dialogue dataset in a university-level CS course, offering insights into how students engage with LLM-chatbot tools in a real learning environment.
3. Dataset
@@
3.1. Participants and Course Context
Our IRB-approved study was conducted among students aged 18 or above, who were enrolled in an AI course at the University of Massachusetts Amherst in the Fall 2024 and Spring 2025 semesters. Across both semesters, 295 students were enrolled in the course, of which 203 agreed to participate in the study. Students who did not participate were also given equivalent access to use the tool over the semester, but their data was not collected. Students were in their 3rd/4th years, had completed their foundational Computer Science course sequences and were taking higher-level elective courses.
@@ 3.1. Participants and Course Context
The programming assignments are Python-based and can be broadly categorized into two types. The first assignment type is a tutorial, where students are given incomplete Jupyter notebooks that guide students to learn fundamental operations in applied machine learning Python libraries, including Pandas, SciKit Learn, and PyTorch. The second assignment type is an applied scenario, where students are given Python starter code that they extend by applying an algorithm or concept taught in class. For example, in assignment 2, students are tasked with creating a text autocomplete engine by implementing various search algorithms to traverse graphs to find word completions. Assignments were altered slightly between semesters, but generally cover the same topics. All assignment instruction details, starter code, and 924 graded submissions are included in the dataset.
Figure 2. The web application used by students in this study. The user interface mirrors the functionality of ChatGPT.
3.2. Data Collection
@@
In order to effectively collect student utterances, we developed and deployed a web application designed to function similarly to consumer LLM chatbot interfaces, shown in Figure 2. As our backing chatbot LLM, we used gpt-4o-mini, the LLM for the free tier of ChatGPT at the time of this study (OpenAI, ). We evaluated this model by testing its capabilities with the libraries and concepts covered in the assignments, and found that the model generally provides accurate guidance on the assignments and implementations that are often correct. Students were instructed that they could use this ChatGPT-based tool, without restrictions, on all assignments in the course. It was emphasized that their approach to using the tool would not affect their grade, and they were encouraged to use the tool in any manner of their choosing. We used the default ”You are a helpful assistant” system prompt, which does not discourage answer-seeking behavior, and did not provide the chatbot with any course materials for reference. Our goal was to encourage students to use our platform instead of seeking an outside application for assignment help, and that the collected interactions would more accurately reflect how students generally use non-instructional, consumer chatbots.
Table 1. Dialogue act schema for categorizing student–LLM interactions, with counts by semester.
Broad Category Specific Dialogue Acts f24 s25 Total
Writing Write Code, Write English, Conversion, Summarize, Other 1664 2672 4336
Editing Edit Code, Edit English, Other 153 265 418
Contextual Questions Assignment Clarification, Code Explanation, Interpret Output, Other 1216 1722 2938
Conceptual Questions Programming Language, Python Library, Computer Science, Programming Tools, Mathematics, Other Concept 2204 2995 5199
Verification Verify Code, Verify Report, Verify Output, Other 293 354 647
Context Assignment Information, Error Message, Code, Other 1191 1837 3028
Off Topic Chit-Chat, Greeting, Gratitude, Other 128 122 250
Misc Other 15 20 35
3.3. Filtering of Student Identifiable Information
@@
To remove any Personally Identifiable Information (PII) from the conversations and assignment submissions, we systematically analyze the conversations to uncover patterns of PII exposure. We first randomly sample 12 conversations from each course topic and read through the conversations, scanning for traces of PII. Our main finding of potentially sensitive student information is in the form of pasted terminal outputs and error messages, which sometimes contain PII, such as a user’s directory or GitHub account name. Compiling these common patterns, we then develop a regular expression-based script to remove these patterns and any direct matches to known user PII in the course roster. Across the entire dataset, this process removes 6,413 cases of possible PII exposure. Then, we sample 12 additional conversations from each topic and performed the same procedure. In our second pass of the second set of conversations, we did not find any additional patterns or cases of PII exposure. We acknowledge that, although seemingly effective, this process does not guarantee the removal of all PII exposures.
3.4. Data Statistics